The blueprints for a patio are shown at the bottom. The patio is a trapezoid. If the cost of the patio is $4.50 per square foot, what will be the total cost of the patio?
base1=18ft
height =15ft
base 2 =24ft

Answer by rfer(16322)

make a triangle
base 6, heigth 15
make a rectangle 18 by 15
find the two areas and add together
--------------
6*15*.5=45 sq ft triangle
--------------------------
18*15=270 sq ft rectangle
-------------------------------
total sq ft
270+45=315 sq ft
-------------------------
315*4.50=$1,417.50